OpenStudy (zab505):

What is 6c3 simplified?

OpenStudy (mathstudent55):

Combinations?

OpenStudy (anonymous):

6c3 written in different form is: \[\left(\begin{matrix}6 \\ 3\end{matrix}\right)\] and formula for that is: \[\left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{ n! }{ k!*(n-k)! }\]

OpenStudy (anonymous):

now you know 6 is n and 3 is k. plug the numbers into the formula

OpenStudy (anonymous):

am going to make an assumption that "6c3" is referring to "Combinations" notation, ₆C₂ , which is the answer to the Q: "How many ways can we choose 2 distinct objects from 6 distinct objects, where order is NOT important?" For example, given the letters A, B, C, D, E, and F, How many different two-letter combinations can be we make? (but where AB is equivalent to BA, so don't count BA as being different from AB) ₆C₂ = 6! ÷ [(6–2)! x 2!] ₆C₂ = 6! ÷ [4! x 2!] ₆C₂ = 720 ÷ [24 x 2] ₆C₂ = 720 ÷ 48 ₆C₂ = 15 ➞ There are 15 ways can we choose 2 distinct objects from 6 distinct objects, where order is NOT important.

OpenStudy (zab505):

Ask yahoo? Knowel

OpenStudy (anonymous):

no yahoo sucks

OpenStudy (anonymous):

my uncle taught me that example